The​ average, or​ mean, D​, of three exam​ grades, w​, r​, and v​, is given by the following formula.
D = w+r+v over 3
​(a) Solve the formula for v.
​(b) Use the formula in part​ (a) to solve this problem. On your first two​ exams, your grades are 86​% and 89​%: w=86 and r=89. What must you get on the third exam to have an average of 86​%?

Answer by Boreal(15235)

D=(w+r+v)/3; 3D=w+r+v; v=3D-w-r
Substitute
v=258-86-89
v=83%